### v1

# Intrinsic Geometry of the Stock Market from Graph Ricci Flow

Abstract
We use the discrete Ollivier-Ricci graph curvature with Ricci flow to examine the intrinsic geometry of financial markets through the empirical correlation graph of the NASDAQ 100 index. Our main result is the development of a technique to perform surgery on the neckpinch singularities that form during the Ricci flow of the empirical graph, using the behavior and the lower bound of curvature of the fully connected graph as a starting point. We construct an algorithm that uses the curvature generated by intrinsic geometric flow of the graph to detect hidden hierarchies, community behavior, and clustering in financial markets despite the underlying challenges posed by a highly connected geometry.
preprint: XXX
The study of the economy as a complex dynamical system has been the focus of a great deal of attention over the past two and a half decades in the field of econophysics[1]. Stock correlations have emerged to become a crucial quantity in the study of stock markets.
The empirical correlation matrix $C_{ij}$ based on the time series of returns $r_{i}$ for stock $i$ is computed from the Pearson correlation coefficients by the usual definition with mean $\bar{r}_{i}$ and variance $\sigma_{i}$: $C_{ij}=\frac{<(r_{i}-\bar{r}_{i})(r_{j}-\bar{r}_{j})>}{\sigma_{i}\sigma_{j}}$
We can then associate a distance that defines an edge weight given by $w_{ij}=\sqrt{2(1-C_{ij})}$ which is our fully connected graph, with all positive weights.
A remarkable property of stock returns is their positive correlation which leads to the emergence of a market index[2]. The spectral properties of the stock correlation matrix have been extensively analyzed and constitute an insightful tool in the risk management of stock portfolios.
On the other hand, the correlation matrix can be used to construct a network with the nodes given by the stocks, which provides a weighted, undirected, fully connected graph.
Existing techniques seek to reduce the connectivity of the graph, by pruning links, to make it more amenable to study. For instance, most standard techniques rely on the minimal spanning tree (MST) to reduce the connectivity of the network[3],[4]. Pruning links to create an MST changes the topology and geometry of the network. In addition, the edge weights of the correlation matrix are very closely spaced, which poses a challenge for MST methods, since their outcome is sensitive to slight differences in the edge weights. Other clustering techniques make assumptions about the geometry of the system or the number of clusters involved (k-means, spectral clustering)[5],[6],[7].
It would be attractive to use recent advances from graph theory, which could shed light on the geometric structure of this network using the full connectivity information without further assumptions. Recent developments in graph theory originating from the adaptation of differential geometry techniques to discrete graphs have shown great potential [8],[9],[10],[11],[12],[13],[14]. The Ricci flow[15], used with great success to solve the Poincaré conjecture[16],[17],[18], is an intrinsic flow of the Ricci curvature tensor, which, coupled with surgery, can be used to decompose a manifold into a connected sum of simpler manifolds. Given the dense connectivity and close distribution of the edge weights of our network, techniques of intrinsic geometric flow provide an attractive alternative to reveal the hidden hierarchy of the data without imposing an external structure.
Figure 1: NASDAQ-100 complete graph
Edge-centric approaches involving graph curvature are increasingly emerging as the method of choice to analyze complex graphs. Combinatorial Forman-Ricci[19],[14],[20] curvature and probabilistic Ollivier-Ricci curvature[8],[9],[10] (ORC) are key developments that have permitted the adaptation of differential geometry methods to discrete structures such as graphs. The correlation matrix, being a fully connected graph, already provides an important indicator of the geometry of the system. Graph curvature has been evoked in the context of the stock market but in these works the network was again replaced by the MST[21],[22],[23],[24]. In this letter, we use the Ollivier-Ricci curvature to construct and analyze the weighted network generated by the stock market as a weighted complete graph. Our main result is to propose a discrete Ricci flow algorithm with surgery based on curvature along neckpinch singularities which allows us to analyze the structure and intrinsic geometry of fully connected graphs with positive curvature, such as the stock market. Detailed comparisons with state-of-the art techniques such as MST and PCA are the subject of a longer paper[25].
Hamilton’s Ricci flow deforms the metric in a fashion analogous to the heat equation,
$\frac{\partial g_{\mu\nu}}{\partial t}=-2R_{\mu\nu}$ (1)
where $g_{\mu\nu}$ is the metric tensor and $R_{\mu\nu}$ the Ricci curvature tensor. This second-order non-linear PDE smooths out irregularities in the metric but leads to singularities that can removed by a procedure known as surgery. There are two steps in adapting Ricci flow to graphs. A breakthrough in adapting this technique to discrete systems came from the definition of coarse curvature or Ollivier-Ricci curvature (ORC) from the work of Ollivier[8],[9],[10]. The second step is devising the flow, as described below.
The Ollivier-Ricci curvature is based on the optimal transport of probability measures associated with a lazy random walk. The discrete Ollivier-Ricci curvature $\kappa_{ij}$ on graph edge $ij\in E$, for a graph $G=(V,E,w)$ with vertices $V$, edges $E$ and weights $w$, compares the Wasserstein distance between two nodes with the shortest distance or edge weight through
$\kappa_{ij}=1-\frac{W(m_{i},m_{j})}{d(i,j)}$ (2)
where $d(i,j)$ is the shortest distance between nodes $i$ and $j$. The Wasserstein distance is defined in the appendix and is related to the probability measures on each site. As explained in the appendix, we use a lazy random walk probability measure, with idleness parameter $\alpha=1/2$. Therefore, we drop the $\alpha$ index, with the understanding that $\alpha=1/2$ in this work.
From this definition, we see that if the nodes $i$ and $j$ belong to the same community, they tend to share neighbors. Therefore the cost of transporting mass is smaller and the Wasserstein distance is no greater than $d(i,j)$. Thus intra-community edges have a positive Ollivier-Ricci curvature. On the other hand, if nodes $i$ and $j$ are from different communities, they tend to have fewer neighbors in common. Therefore, transporting mass from $i$ to $j$ necessarily involves passing through the edge $ij$ and the Wasserstein distance is larger than $d(i,j)$. Therefore the curvature of edge $ij$ is negative and the link is hyperbolic. On a flat graph, the Wasserstein and $d(i,j)$ distances are identical and the curvature is zero. If all the links in a graph have zero curvature, then the graph is called Ricci-flat. The regular grid is an example of a Ricci-flat graph.
With this definition of coarse curvature, we can now build the Ricci flow, to generate a time-dependent family of weighted graphs $G(V,E,w(t))$, such that the weight $w_{ij}(t)$ on edge $ij$ changes proportionally to the Ollivier-Ricci curvature $\kappa_{ij}(t)$ at time $t$. Ollivier[9],[10] suggested a flow with continuous time $t$,
$\frac{d}{dt}{w_{e}(t)}=-\kappa_{e}(t)w_{e}(t)$ (3)
where edge $e\in E$, with Ollivier-Ricci curvature $\kappa_{e}$ and edge weight $w_{e}(t)$. Jost and co-workers used a discrete time algorithm with the Forman-Ricci curvature. Sia and co-workers[26] introduced an algorithm based on surgery with curvature but links were not updated. Ni and co-workers[27] suggested a discrete-time algorithm while updating the edge-weights. Each iteration of this algorithm consists in updating all edge weights of the network simultaneously. The $l^{th}$ step of this process gives
$w_{ij}^{(l+1)}=(1-\kappa_{ij}^{(l)})d^{(l)}(i,j)$ (4)
where $w_{ij}^{(l)}$ is the weight of edge $ij$ at the $l^{th}$ iteration, $\kappa_{ij}^{(l)}$ is the Ricci curvature of edge $ij$ at the $l^{th}$ iteration, and $d^{(l)}(i,j)$ is the shortest path distance induced by the weights $w_{ij}^{(l)}$. Initially $w_{ij}^{(0)}=w_{ij}$ and $d^{(0)}(i,j)=d(i,j)$, the initial weights and distances of the graph. Therefore at the $l^{th}$ iteration, the edge weights are replaced by the Wasserstein distance to define the graph at the $(l+1)^{t}h$ iteration. The procedure is then repeated to obtain the new Ricci curvature. The Ricci flow has the effect of bringing the positive edges closer and pushing the negatively curved edges further apart. This effect is used to perform surgery on the network and separate the network into communities.
With this definition of the Ricci curvature, propagating these values by the Ricci flow tends to exaggerate the differences between the different types of links. Links with positive curvature tend to cluster together and links of negative curvature tend to move further and further part under the Ricci flow, while flat links are unaffected. While community detection in graphs is a very active field of research[28],[29],[30], this brings a new approach to cluster complex networks[20],[27].The multiscale structure of networks has been studied using dynamical ORC[31].
Figure 2: Illustration of Wasserstein distance and Graph curvature on fully connected and tree-like graphs
How do we look for hierarchies in such a network? In our work, we consider as our starting point the fully-connected graph with equal weights. In the literature, the fully connected graph with equal edge weights is often referred to as the discrete equivalent of the spherical manifold. Therefore, the expectation of communities bound together and separated by well-traveled or hyperbolic links does not hold here. The eigenvalues of the Laplacian of the complete graph are known to be 0 and $\frac{n}{n-1}$ with multiplicity n1n-1. If we consider the complete graph with N nodes and $\alpha=1/2$ the Wasserstein calculation is illustrated in Fig. (2) for a 5 node graph. We see that $W_{ij}\leq 1$ therefore, the Ricci curvature is always positive. So for a fully connected graph with N nodes and all equal edge weights, the Ricci curvatures of all edges are equal and given by $\kappa_{ij}=(1-\alpha)\frac{n}{n-1}$= $\kappa$. It is known[12] that the Lin-Lu-Yau-Ricci curvature $\lim\limits_{\alpha\rightarrow 1}\frac{\kappa_{ij}}{(1-\alpha)}=\frac{n}{n-1}$ of the complete graph $K_{n}$ is the only graph with a constant curvature greater than 1, which fits with the value above for $\alpha=1/2$. The eigenvalues of the Laplacian of the complete graph are the same as the Lin-Lu-Yau-Ricci curvatures. In the case of the NASDAQ 100 network, if we approximated equal edges, with $\alpha=1/2$, we would have $\kappa=0.505$. Thus, this complete graph is a highly degenerate graph. It is expected that when the edge weights are not all equal, this degeneracy will be broken.
In Fig. 1 we present our construction of the NASDAQ-100 graph with empirical edge weights, constructed from real-life data. We recovered the values of stock prices from Yahoo! Finance[32] using the components of the NASDAQ-100 Index from [33]. We calculate the returns and average over a period of five years from 1/1/2020-20/11/2024. Based on the edge weights, we obtain a histogram of Ricci curvatures[27],[34]. The histogram is shifted downward from the complete graph at 0.505, though being all positive retain the spherical aspect of this graph. This is reflected in the downward shift of the distances from 1. However, the curvatures are very closely distributed around the median, and a priori, this is not a very favorable situation to select a given link for surgery. The correlation of this graph is $38\%$, which means that there is an average edge-weight of 0.78. In Fig. 1 we can see the distributions of the edge weights and they are shifted downwards from 1 and are all very tightly distributed around a median value of 0.8, in accordance with the correlation. Here the $\kappa_{ij}$ are all positive which corresponds to a graph of positive curvature, while the edge weights are not identical in a fully connected graph, due to the connectivity. With the Ricci coarse curvature, we are able to use the full correlation matrix to build a network and carry our Ricci flow without any assumptions about the number of clusters.
The geometry of three and four manifolds under Ricci flow was studied extensively by Hamilton. A limited version of the Poincaré theorem for Riemannian manifolds of strictly positive curvature was proved by Hamilton[15]. In the case of manifolds with positive curvature, it is known that evolution under Ricci flow leads to neckpinch singularities which are removed by surgery[35]. A dumbbell-like manifold under Ricci flow develops a neckpinch singularity and surgery is carried out along the cylindrical link[36].
First, let us consider the case of a Ricci flow on the fully connected graph with equal edge weights. It is clear that if we look at the graph with equal weights, if we do not update edge weights using the algorithm, the Wasserstein distance and Ricci curvature are unchanged. If we update the edge weight, Ricci flowing the graph will lead to shortening the edge weights and eventually to a singularity. Yet, the graph remains self-similar, in the sense that it is a fully connected graph with equal edge weights, which decrease with time, eventually leading to a point-like singularity. All edges remain equal, and there is no question of separating this graph into communities through different edges. Other approaches exist in the literature where the normalized Ricci flow is used, in which case there is no singularity, though self-similarity remains.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 3: NASDAQ-100 Ricci flow after 5, 10 and 20 iterations with surgery on the 20th iteration
The situation is entirely different when we consider the fully connected NASDAQ-100 graph with distinct empirical edge weights obtained from the stock market. The challenge here is to analyze the empirical NASDAQ graph to highlight features and communities. In the continuum the neckpinch singularities are identified, and surgery is performed to separate the original manifold into a connected sum of manifolds while avoiding collapse. In the discrete context, the key step is to be able to identify the neckpinch singularities. Here we will attempt a similar decomposition of our graph with the expectation that we will find spherical ‘communities‘. These are not connected by hyperbolic links, therefore the challenge is to isolate them. One was to do this would be the development of a locally flat neckpinch singularity, which can be embedded into a locally flat manifold.
We expect that Ricci flow should shrink the edges of a community to regions of large positive curvature. This is indeed the case. If we use edge weights as a measure, it is still not clear how to carry out surgery in the fully connected manifold using distances. In [26] edge curvature alone without Ricci flow was used, with the argument that triangles are a good index of community formation. We suggest in this work, based on Ricci flow and the intuition obtained from the fully connected graph, that curvature is a good metric to use.
We carried out the Ricci flow on the NASDAQ graph with $\alpha=\frac{1}{2}$ and a discrete time step[34]. In Fig. (3) we see that after just 5 iterations of Ricci flow, the distribution of edge weights begins to flatten out. The distribution of curvatures is more remarkable; we begin to see the data split into two peaks. This confirms our intuition that edge curvature is a sensitive measure of clustering in data. We expect that a neckpinch singularity develops.
From Fig. (3), we notice that after just 5 iterations of Ricci flow, the histogram of edge weights has become much wider. The curvature histogram is even more pronounced, with the appearance of two peaks. We make a scatter plot of edge weight vs. curvature, and this is widely dispersed. Visually we see that fully connected network has already started separating into two communities. After 10 iterations, these features are much more pronounced, with the appearance of a dumbbell-shape and a neckpinch between the two communities. From the spread of the scatter plot, it is not yet clear how to perform surgery. By 20 iterations, the histograms of edge weight and edge curvature are completely changed. The system is split into two communities. The scatter plot provides us with the means to select the links for surgery. We perform surgery along the zero-curvature links, which correspond to the long flat links that separate the two communities.
After 20 iterations, we see a clear separation between different clusters. This is better seen in a scatter plot. We plot the edge weight and curvature scatter plot. We find the links flattening and the curvature approaches zero.
From Lin and Yau[13],[37], the infinite weighted graph has a lower bound to the ORC given by $\kappa_{ij}\geq\kappa_{ij,min}=\frac{2}{d_{i}}+\frac{2}{d_{j}}-2$, where $d_{i}$ is the degree of link $i$. The lower bound has been further developed for hypergraphs[38]. In the limit of the infinite complete graph with equal edge weights, this gives a value for $\kappa_{ij,min}=\frac{2(3-n)}{(n-1)}$. For the NASDAQ-100 graph this gives $\kappa_{ij,min}\approx-1.96$. Naively, this would correspond to a length of $d_{ij}\approx 2.96$ as the longest edge length possible. We find after 20 iterations that the graph splits into two communities, with very short edge weights and positive curvature. These two communities are separated by long, flat links with curvature zero and edge weights around 3. It is interesting to note that the fully connected graph which satisfies strict positivity of the curvature before the Ricci flow, still obeys the lower bound condition, even with unequal edge weights. This brings us to the question of surgery in such a graph, it would seems appropriate to use the flattening that develops along the neck pinch singularity to perform surgery along the zero-curvature links. This suggests that the geometry of the NASDAQ-100 graph is a prominent influence on its properties. The reason is that the degree of the graph seems to play a key role in influencing the lower bound, even without equal edge weights.
With this procedure for surgery, based on curvature, we can analyze the system through several levels of Ricci flow followed by surgery. In this work, we sketch the clustering obtained by our technique, with a more detailed financial analysis of the clusters presented in another work[39]. The NASDAQ-100 is a tech-heavy index. After the first level of Ricci flow, followed by surgery using the method described above, the cluster splits into two communities of 66 and 35 stocks. The larger cluster is clearly tech-driven and we first focus on this. The second level of Ricci flow leads to a basket of 16 semiconductor stocks. It is encouraging to see that Ricci flow can identify the semiconductor stocks. However Nvidia is absent from this cluster. The level 3 Ricci flow of the 50 stock cluster leads to the separation of a cluster of 10 stocks. When we examine these stocks, it is clear that they are outliers. Moderna, which was a very poor performer was subsequently dropped from the NASDAQ in December 2024. Analyzing the 40 stock cluster leads to the separation of biotech stocks and Supermicro stock. The cluster of the remaining 35 stocks splits into two roughly equal groups of 17 and 18 stocks. We name the 17 stock cluster ’Big Tech’ which contains Nvidia and Amazon amongst others, so Nvidia finds its position in the market makers stock over the more limited semiconductor stocks. Tesla is in the other cluster.
Figure 4: Level 5 clustering of NASDAQ-100
The other level 1 cluster of 35 stocks seems to contain non-cyclicals, which we are then able to cluster into Healthcare staples, Energy staples and Industrials. See Fig. (4). The detailed cluster components are listed in the tables in the appendix.
A detailed analysis of the clusters from the point of view of econophysics is underway[39]. We are carrying out more systematic comparisons with MST and variations of $\alpha$, which will be presented in a longer paper[25]. It is also important to delve deeper into the clusters provided by this method and to study a larger, less tech-heavy index such as the S&P500 index. The behavior around crashes is also being investigated. .
In conclusion, we study the NASDAQ-100 stock index from the point of view of a fully connected network and provide an algorithm to carry out surgery based on curvature values after Ricci flow on this graph. We are able to develop a novel clustering technique on this fully connected graph, without assumptions about the number of clusters. We are able to identify outliers. Our curvature based approach provides an attractive alternative to MST to study the geometric properties of the stock market. Our technique provides a novel, curvature based criterion to perform surgery on highly-connected graphs with initial positive curvature.
Acknowledgements.
We wish to acknowledge the support of the CNRS and LPTMS.